What is dy/dx called?

Grade:7

8 Answers

aditya
9 years ago
its a differential..... differentiating y with respect to x
mohini mohan behera
9 years ago
slope of the tangent to curve y=f(x).
ng29
9 years ago
let us given a function y=f(x)
dy/dx represents the differentiation of function y=f(x) with respect to x 
eg .
given  y=x^2
then dy/dx=2x
approve if useful
suman gupta
8 years ago
DY/DX means that we can simply tell that diffentiation means process of sampling and by example we can tell that                           dy/dx= 2 it means that x is 1 unit and y is 2 unit…..
                                                                                                                        thank u......
 
Prajwal Kavad
8 years ago
it is called differentiation
 
Adarsh
8 years ago
It means derivative of y with respect to x.
 
Hope you understand my answer.
Shreyas Patil
7 years ago
de
Priyansh
6 years ago
it is a symbol used for differentiation called  derivative of x with respect to y...........................................................................